OpenStudy (clarence):

"Assume that the graph of the polynomial p(x)=ax2+bx+cpasses through the point (−1,2)and has a horizontal tangent at (12,−9). Find the sum a+b+c." I've already substitute (-1, 2) into the polynomial equation to get a - b + c = 2 and differentiate the original polynomial to get 2ax+b=0 and replaced x with 12 to get 24a + b = 0 as well as utilizing (12, -9) to get 144a + 12 b + c = -9 but I have no idea where to go from there..

OpenStudy (asnaseer):

your steps are correct so far

OpenStudy (asnaseer):

you now have 3 equations with 3 unknowns

OpenStudy (asnaseer):

solve these to find a, b and c and then calculate a+b+c

OpenStudy (clarence):

I've tried to use row operations but was unsuccessful, is there a particular order that I should place these equations?

OpenStudy (asnaseer):

You could just use simple substitution, e.g. start by eliminating c using the last equation

OpenStudy (clarence):

I'm not quite sure I understand you, use 144a + 12b + c = -9 to eliminate c?

OpenStudy (asnaseer):

c = -9 -12b -144a from that equation now substitute this into your first equation

OpenStudy (clarence):

Oh, the a-b+c=2 equation?

OpenStudy (asnaseer):

yes :)

OpenStudy (asnaseer):

you will then be left with 2 equations in two unknowns - a and b

OpenStudy (clarence):

How does that help? I have -13b-143a=11 right now.

OpenStudy (asnaseer):

and you also have: 24a + b = 0

OpenStudy (asnaseer):

which gives you: b = -24a substitute that into: -13b-143a=11 and solve for a

OpenStudy (anonymous):

ok what do you need help with, solving for a?

OpenStudy (clarence):

a would equal 2/169?

OpenStudy (asnaseer):

not quite - you must have made a simple algebraic mistake

OpenStudy (clarence):

Did you not get 312a-143a=2?

OpenStudy (anonymous):

-13b-143a=11 -13(-24a) - 143 a = 11 312a - 143a = 11 169a = 11 /169 on both sides a = 11/169

OpenStudy (asnaseer):

remember the negative sign

OpenStudy (anonymous):

me? It went away

OpenStudy (asnaseer):

I am talking to @Clarence

OpenStudy (anonymous):

ok sorry just trying to help

OpenStudy (clarence):

So b = - 264/169?

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

good job!!

OpenStudy (asnaseer):

now you are just left with c to calculate and then you can work out what a+b+c equals

OpenStudy (clarence):

c being 63/169?

OpenStudy (asnaseer):

yup :)

OpenStudy (anonymous):

wow sounds like you got it down, now just add them together

OpenStudy (clarence):

So all together it'd be -190/169?

OpenStudy (asnaseer):

yes - well done! :)

OpenStudy (anonymous):

correct awesome job

OpenStudy (clarence):

Oh I get it now, thanks!

OpenStudy (asnaseer):

why do I feel like there is an echo in here :)

OpenStudy (clarence):

Probably because there is ;) Thanks again guys/girls!

OpenStudy (anonymous):

haha you keep typing and replying right before i do and i don't see it sorry

OpenStudy (clarence):

Aha, all good :)